In some aquatic ecosystems, nitrate (NO3–) is converted to nitrite (NO2–), which then decomposes to nitrogen and water. As an example of this second reaction, consider the decomposition of ammonium nitrite: NH4NO2(aq) = N2(q)+2H2O(l)
What would be the change in pressure in a sealed 10.0 L vessel due to the formation of N2 gas when the ammonium nitrite in 1.40 L of 0.900 M NH4NO2 decomposes at 25.0°C? answer in atm
1
Expert's answer
2019-12-09T08:05:47-0500
"NH_4NO_2 \rightarrow N_2 + 2H_2O"

Find moles of NH4NO2

"n=c\times V= 0.900\times 1.40 = 1.26 mol"


Accordring to equation "n(N_2) = n(NH_4NO_3) = 1.26 mol"


Use Ideal gas law to find pressure:


"PV = nRT"

"T(K) = 273 + 25 = 298 K"

"V= V_{vessel}-V_{solution}= 10.0-1.4 =8.6 L"



"P\times 8.6 = 1.26\times 0.082\times298"

"P = 3.58 atm"